When copper is placed in silver nitrate a single displacement reaction occurs, producing solid silver as one of the products. When a group of students carried out this reaction they calculated that 2.50 g of solid silver should be produced given the amounts of the reagents they used. However, they were only able to isolate 2.13 g of silver. What is the percent yield for this reaction?


1
Expert's answer
2021-12-20T00:50:01-0500


2Ag+S→Ag2​SThe atomic mass of Ag and S are 108 g/mol and 32 g/mol respectively. The Molar mass of Ag2​S=2(108)+32=248 g/mol.2 moles of Ag gives 1 mole of Ag2​S.27 g of Ag will give 108×2
27​×248=31 of Ag2​S.Hence, the theoretical yield is 31 g.However only 25 g of silver sulfide was collected.Hence, the actual yield is 25 g.The percentage yield is 31
25×100​=81 %.